OpenStudy (anonymous):

I need to understand this problem. Integrate 7xcos^4x^2 dx

OpenStudy (anonymous):

\[\int\limits_{}^{}7xcos ^{4}x ^{2} dx\]

OpenStudy (anonymous):

Allright. First we need to take care of that x^2 inside the cosine.

OpenStudy (anonymous):

Do you see how we would do that?

OpenStudy (anonymous):

By using a u sub?

OpenStudy (anonymous):

Good job :) .

OpenStudy (anonymous):

So u=x^2 and du=2x dx

OpenStudy (anonymous):

or du/2 = x dx

OpenStudy (anonymous):

Recall that 7 is a constant which can be pulled outside the intergal.

OpenStudy (anonymous):

So what do you obtain?

zepdrix (zepdrix):

After you get it into the `u form` like you had in the last post, I would finish it up by using the `Cosine Reduction Formula` like we did earlier. :) If applying double angle over and over works better for you though, you can do that.

OpenStudy (anonymous):

okay, awesome. The 7 is part of what was throwing me off. So now we have \[7\int\limits_{}^{} \cos ^{4}(u) \frac{ du }{ 2 }\]

OpenStudy (anonymous):

Yep that's what I am doing @zepdrix :P . Just walking him through :) .

OpenStudy (anonymous):

her* :P

OpenStudy (anonymous):

du/2 is really 1/2 * du . We can also pull out that 1/2 in front of the integral :) .

OpenStudy (anonymous):

Sorry X) .

OpenStudy (anonymous):

so we CAN use the cosine reduction formula?

OpenStudy (anonymous):

Yep :) .

OpenStudy (anonymous):

We obtain: